# Towards Tight Convex Relaxations for Contact-Rich Manipulation

### v1

@@ VIII Experiments
Finally, we demonstrate the feasibility of the obtained motion plans on a Kuka LBR iiwa 7 R800 7-DOF robotic arm, with a T-shaped slider object. A picture of the experimental setup is shown in Figure 1. We let the integration constant in (11) be $c=0.3$. We find that the plans often perform well open-loop. For extra stability, we use a feedback controller to execute the plans on hardware, and employ a hybrid Model-Predictive Controller (MPC)commonly used for pushing tasks [49, 42]. A selection of qualitatively different plans and the hardware execution can be seen in Figure 5.
@@ IX Conclusion and Future Work
In this work, we present a framework for planning near-globally optimal trajectories for contact-rich systems. We demonstrate how the dynamics of a contact-rich system can be described as a QCQP, and leverage semidefinite relaxations to approximate the planning problem in a fixed contact mode as a convex program. This convex approximation of the dynamics enables us to plan complex trajectories over mode sequences by leveraging the SPP in GCS framework from [1]. The entire planning problem can therefore be solved as a single convex optimization followed by a cheap rounding strategy based on local optimization. As an initial application, we apply our framework to the planar pushing problem and demonstrate that our method empirically generates near-globally optimal solutions. We are currently working on applying our method to more complex contact tasks.
We are additionally exploring how further techniques from the literature on SDP relaxations can be used to automatically provide even tighter relaxations in a reasonable computational budget. Additionally, while the proposed approach is more efficient than prior global optimization methods, the planning times are still too slow for fast, reactive planning. We are currently working on better exploiting our highly structured problem formulation, both through the use of a custom solver and more efficient construction of the optimization problem. Additionally, there is a rich literature on both exploiting special structure particularly in SDP s [66, 69], as well as approximating SDP s using simpler cones [70, 71] which could potentially reduce solve times by several orders of magnitude. Future work will explore the ability of these reduction methods to accelerate the planning.

### v2

@@ VIII-C Execution on real hardware
Finally, we demonstrate the feasibility of the obtained motion plans on a Kuka LBR iiwa 7 R800 7-DOF robotic arm, with a T-shaped slider object. A picture of the experimental setup is shown in Figure 1. We experimentally determine approximate values for the friction parameters from a handful of open-loop executions, and find the friction coefficient and integration constant for the contact between the slider and table in (7) to be $\mu_{S}=0.5$ and $c=0.3$, and the friction coefficient between the pusher and slider in (10) to be $\mu=0.05$. We find that the plans generated with these friction parameters often perform well open-loop. For extra stability, we use a feedback controller to execute the plans on hardware, and employ a hybrid Model-Predictive Controller (MPC)commonly used for pushing tasks [42, 51]. Two different plans and the corresponding hardware execution can be seen in Figure 5, and a collection of hardware demonstrations can be seen in the video at https://bernhardgraesdal.com/rss24-towards-tight-convex-relaxations/.
@@ IX Conclusion and Future Work
In this work, we present a framework for planning near-globally optimal trajectories for contact-rich systems. We demonstrate how the dynamics of a contact-rich system can be described as a QCQP, and leverage semidefinite relaxations to approximate the planning problem in a fixed contact mode as a convex program. This convex approximation of the dynamics enables us to plan complex trajectories over mode sequences by leveraging the SPP in GCS framework from [1]. The entire planning problem can therefore be solved as a single convex optimization followed by a cheap rounding strategy based on local optimization. As an initial application, we study our framework on the planar pushing problem and demonstrate that our method empirically generates near-globally optimal solutions. We are currently working on applying our method to more complex contact tasks.
While the proposed approach is more efficient than prior global optimization methods, the planning times are still too slow for fast, reactive planning. We are currently working on better exploiting our highly structured problem formulation, both through the use of a custom solver, more compact problem formulations, and more efficient construction of the underlying optimization problem. Additionally, there is a rich literature on both exploiting special structure particularly in SDP s [69, 71], as well as approximating SDP s using simpler cones [72, 73] which could potentially reduce solve times by several orders of magnitude. Future work will explore the ability of these reduction methods to accelerate the planning. In addition, we are exploring how further techniques from the literature on SDP relaxations can be used to automatically provide even tighter relaxations in a reasonable computational budget.